# Fenglin Liu1, Jinge Wu211footnotemark: 1, Hongjian Zhou1, Xiao Gu1, Soheila Molaei1,&Anshul Thakur1, Lei Clifton3, Honghan Wu2,4, David A. Clifton1,5

### v1

@@ Results
We first build a benchmark MedRisk for evaluating LLM’s performance for risk prediction. To build an accurate benchmark dataset consisting of diverse risk problems, we first collect all available evidence-based tools, i.e., clinical calculators, from clinical-standard source MDCalc [41]. We then review and exclude tools that are not directly predicting disease risks (e.g., those predicting Pregnancy Due Dates and BMI), retaining 387 tools that predict 154 diseases across different scenarios. We then use the APIs provided by MDCalc [41] to generate over 15,000 questions with free-text reference answers by requesting GPT-4o [2] in Azure to randomly select appropriate tool parameters, approximately 20 quantitative and 20 qualitative questions for each tool. Next, we manually identify and exclude data samples that are unrelated to risk prediction, obtaining 12,352 risk questions across 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems. For robust and efficient evaluation, we convert these questions and free-text references to multiple-choice format by combining the correct answer with three wrong choices. Finally, we divide the benchmark into MedRisk-Qualitative and MedRisk-Quantitative, with 6,176 data samples each, according to their question and answer types. For example, the question ‘What is the estimated risk of postoperative pulmonary complications?’ (Answer: High risk) belongs to Qualitative, while the question ‘What is the Risk Score for Asthma Exacerbation (RSE) for this patient?’ (Answer: 30-37%) belongs to Quantitative. Figure 1(b) shows the statistics of the top-30 diseases, top-30 symptoms, top-15 specialties, and top-15 organ systems in the MedRisk benchmark.
To comprehensively evaluate LLMs’ performance for risk prediction, we collect 13 different types of representative LLMs and agents. 1) Public LLMs: We collect multiple representative open-source public LLMs across different numbers of model parameters, i.e., Mistral-7B [34], LLaMA-3-8B [35], Mixtral-8x7B [36], and LLaMA-3-70B [35]; We also collect two public medical LLMs: PMC-LLaMA-13B [11] and Meditron-70B [15]. They are developed by fine-tuning general LLMs on a large corpus of medical datasets [6] to learn rich medical knowledge , achieving improved performance in medical tasks. 2) Commercial LLMs: We further collect the currently most powerful commercial LLMs, GPT-4o-2024-08-06 [3], o3-mini-2025-01-31 [16], o1-2024-12-17 [4], and GPT-4.5 [37]. 3) AI Agents: Recently, AI agents [43, 44], which assign different LLMs to different roles to collaboratively perform tasks, have demonstrated better task performance than single LLMs in traditional domains such as gaming and programming. Therefore, for a comprehensive comparison, we further collect three well-known and recent AI agents, i.e., ReAct [38], BOLAA [39], and Chameleon [40]. We adopt LLaMA-3-8B [35] as the backbone of our system to re-implement them for a fair comparison.
Main Evaluation
In Table 1, we report different methods’ performance on the most common cases across disease, symptom, specialty, and organ system and their overall performance. RiskAgent not only outperforms different types of LLM in all cases, respectively, but also achieves the highest overall accuracy of 78.34% and 76.33% in MedRisk-Qualitative and MedRisk-Quantitative, respectively. It indicates the robustness of our method across different diseases, symptoms, specialties, organ system, and question types, providing a more comprehensive risk prediction solution than previous methods. In detail, compared with large commercial LLMs (e.g., GPT-4o, GPT-o1, and GPT-o3-mini) and open-source LLMs (e.g., Meditron-70B and LLaMA-3-70B), our RiskAgent outperforms them by $>$20.0% accuracy, with much (at least 10 times) fewer parameters. This is desirable in resource-limited medical settings. Our method also achieves the best results among AI agents, with accuracy surpassing previous methods by up to 35.83% overall accuracy.
To clearly understand the effectiveness of our approach, we visualize the results on MedRisk-Quantitative in Figure 1(c), which compares our method with the competitive LLMs. It shows that our RiskAgent outperforms previous methods in all cases. These encouraging results demonstrate that our RiskAgent achieves more precise and robust decision support than existing advanced LLMs for complex medical tasks. Importantly, our method uses evidence-based tools for prediction and traces the information source (i.e., evidence) behind our system’s decisions, making the results more transparent and reliable than existing approaches. In Figure 3, we further plot the results of five runs for RiskAgent-8B, GPT-4o, and o3-mini. We can see that RiskAgent-8B not only consistently achieves the best results across all cases in different runs but also achieves lower perturbations (STD) than o3-mini in most cases, especially in risk prediction for cardiac.
Figure 4: The robustness of our method: We evaluate the performance of models on five rare diseases (left) and six types of cancer (right). SM: Systemic Mastocytosis; PM: Primary Myelofibrosis; CML: Chronic Myelogenous Leukemia; IPF: Idiopathic Pulmonary Fibrosis.
Evaluation of Rare Diseases and Cancer
In this section, to evaluate our method’s robustness in solving diverse complex diseases, we further conduct experiments on rare diseases and cancer.
For rare (but important) diseases, we follow two clinical-standard rare disease databases, i.e., European Orphanet [^1] https://www.orpha.net/ and the US National Organization for Rare Disorders (NORD) [^2] https://rarediseases.org/, to choose five rare diseases for evaluation. The left part of Figure 4 illustrates the results of our RiskAgent and previous strong LLMs. As we can see, RiskAgent achieves the best results in all cases. In particular, in Amyloidosis, RiskAgent achieves greater improvements than the most recent commercial LLM, GPT-4.5. For cancer, as shown in the right part of Figure 4, we evaluate different methods on six cancers. We can clearly see that our method consistently achieves competitive results with different For example, on Gastric Cancer and Breast Cancer, RiskAgent surpasses previous LLMs by at least 37.5% and 28.57% accuracy, respectively.
Overall, our RiskAgent, with only 8B model parameters, achieves encouraging results across diverse rare diseases and cancers. It demonstrates the robustness of our method and shows the potential of our solution to provide more accurate and robust risk prediction for complex diseases, especially for rare diseases, compared to existing state-of-the-art LLMs.
Figure 5: The generalization ability of our method: We report the overall accuracy of the basic LLMs (blue bars) and the basic LLMs enhanced using our method (red bars). We evaluate a. different variants of the GPT-4o-series LLMs developed at different times and b. LLaMA-3-series LLMs with varying numbers of model parameters. The polyline and the right y-axis show the improvements in different variants. We can see that the more advanced (a) and the larger (b) the basic LLM, the greater the improvements.
Generalization Analysis
In this work, we provide a solution that allows LLMs to collaborate with evidence-based medical tools to produce accurate, reliable medical results. Thus, our method is agnostic to the underlying LLM, making it adaptable to various LLMs and helping them produce accurate and evidence-based medical outputs. To this end, we further explore our method’s generalizability and applicability as LLMs continue to evolve rapidly over time and as the number of model parameters increases, especially considering the rapid development of LLMs and the trend toward increasing model sizes. During evaluation, we select the advanced commercial GPT-4o-series developed at different times (gpt-4o-2024-05-13, gpt-4o-2024-08-06, and gpt-4o-2024-11-20) and the representative public LLaMA-series (LLaMA-3-1B, LLaMA-3-3B, and LLaMA-3-8B), which share the model architecture but differ in model size.
We compare the performance of (i) the basic LLM and (ii) the basic LLM enhanced using our method. Figure 5 shows that our method provides significant performance improvements for all basic LLMs (accuracy of 17.04$\sim$44.27%) on both MedRisk-Qualitative and MedRisk-Quantitative benchmarks. In detail, results on the GPT-4o-series indicate that as models evolve, the performance improvement brought by our method increases. This indicates that if downstream applications do not involve strict privacy regulations, our method can significantly improve the performance of large advanced LLMs, enabling them to accurately solve complex medical problems. Even when privacy concerns are involved, our method can still significantly boost the results of LLaMA-3-series public LLMs across different model parameters. For example, our method improves accuracy by 19.57$\sim$30.52for the 1B and 3B LLMs, enabling them, with only 1B and 3B model parameters, to outperform the commercial GPT-4o basic LLMs. Such encouraging improvements in lightweight LLMs are useful in resource-limited medical institutions.
Meanwhile, we notice that as model sizes increase, the improvements brought by our method become greater. Overall, the improved results across different GPT-4o-series and LLaMA-3-series LLMs show the strong potential of our approach, enabling large and advanced LLMs to deliver accurate medical data analysis and evidence-based predictions. As LLMs continue to evolve rapidly, our method can further enhance their effectiveness in complex, critical medical tasks.
External Evaluation
We further evaluate our method’s generalization ability on an external evidence-based diagnosis benchmark (MEDCALC-BENCH [45]), which requires the models to use medical calculations specifically for clinical diagnosis to perform evidence-based decision support. Note that our RiskAgent is designed specifically for risk prediction; thus, this benchmark evaluates our methods’ generalization ability to new tasks and domains.
For evaluation, we follow previous work [45] to employ zero-shot chain-of-thought (CoT) prompting to report the results. We adopt different backbones, i.e., LLaMA-3-8B [35] and GPT-4o [3], to implement two variants of our method: RiskAgent-8B and RiskAgent-GPT-4o. Table 2 demonstrates the outstanding performance of our approach on external evaluation using the evidence-based clinical benchmark MEDCALC-BENCH. The results clearly show that: (i) The RiskAgent-8B model, with only 8B parameters, surpasses all public LLMs, including the larger LLMs Meditron-70B [15] and Mixtral-8x7B [36] in major cases. (ii) The RiskAgent-GPT-4o model significantly outperforms all baseline models with an overall accuracy of 67.71%, representing a substantial improvement over the standard GPT-4 and GPT-4o. Most notably, our model exhibits exceptional performance in rule-based diagnosis tasks, achieving 61.67% accuracy in the Diagnosis category - more than double the performance of all baseline methods. Meanwhile, the performance improvements are particularly striking in calculation-heavy categories, with RiskAgent-GPT-4o achieving 97.08% on Physical assessments and 95.00% on Dosage calculations - areas where precise quantitative reasoning is critical for patient safety. It can also be verified by our RiskAgent-8B model, which, surprisingly, outperforms all public and commercial LLMs in dosage calculations. These encouraging results demonstrate that our RiskAgent framework not only excels in risk prediction, but also shows comfortable generalization capabilities across diverse diagnostic tasks in previously unseen clinical scenarios. The significant performance gains across both equation-based and rule-based calculations further underscore the robustness of our approach and its potential to provide reliable decision support across various medical contexts.
@@
Public LLMs PMC-LLaMA-13B [11] 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Meditron-70B [15] 0.00 0.00 3.33 0.00 0.00 0.00 3.33 0.38
Mistral-7B [34] 10.09 14.58 1.67 0.00 9.58 7.50 25.00 10.79
LLaMA-3-8B [35] 16.51 25.00 1.67 7.50 11.25 13.75 26.67 16.43
Mixtral-8x7B [36] 22.63 40.00 6.67 17.50 11.25 21.25 15.00 22.35
@@
Commercial LLMs GPT-3.5 [46] 20.49 45.00 11.67 17.50 13.33 10.00 31.67 23.69
GPT-4 [2] 26.30 71.25 48.33 40.00 27.50 15.00 28.33 37.92
GPT-4o [3] 40.98 82.50 36.67 47.50 32.50 22.50 23.33 46.13
@@
Table 2: Evaluation of our RiskAgent, designed for risk prediction, on an external evidence-based clinical diagnosis benchmark (MEDCALC-BENCH[45]), which is categorized into Equation-based (Lab, Physical, Date, Dosage) and Rule-based (Risk, Severity, Diagnosis) diagnosis. All values are reported in percentage (%).
@@
Table 3: Ablation study of our RiskAgent, which includes four main components: Tool Library, Decider, Executor, and Reviewer. We take LLaMA-3-8B [35] as the basic LLM and use accuracy to report our results. We add the results of a public LLM (LLaMA-3-70B [35]) and a commercial LLM (GPT-4o [3]) for comparison.
Ablation Study
In this section, we provide an ablation study in Table 3 to better understand the contributions of each component. Our method introduces four main components, i.e., the Tool Library in the environment, the Decider, the Executor, and the Reviewer, to achieve superior performance.
@@
(i) The results show that all components contribute to improved performance, proving the effectiveness of each component. (ii) By comparing Basic LLM and Setting (a), which introduces the tool library to enable the basic LLM to use the evidence-based tool for risk prediction, we observe that Setting (a) significantly boosts the performance of the basic LLM by more than 14% accuracy, surpassing larger models like LLaMA-3-70B and GPT-4o. This proves the motivation and effectiveness of leveraging evidence-based tools to achieve improved performance on complex medical tasks. (iii) Considering that collaborating with tools involves both tool selection and parameter parsing for tool execution, we further report their accuracy. Comparing Setting (a) and Setting (b), we observe that introducing LLM agents can significantly boost performance—improving tool selection accuracy by 9.66%, parameter parsing accuracy by 30.20%, and overall task performance by 18.99%. This proves the effectiveness of our method in separating the collaboration process into tool selection and parameter parsing through the designed Decider and Executor. By distributing responsibilities, our approach reduces the model’s learning burden, enabling accurate and efficient collaboration with medical tools. (iv) Finally, we notice that the Reviewer further improves overall performance, showing the effectiveness of reflecting on the analysis process and thereby correcting potential errors. Such a reflection process is very similar to the ‘Aha moment’ in the recently popular DeepSeek-R1 model [47], further proving the effectiveness of our method.
Overall, the ablation study shows that both collaboration with evidence-based medical tools and the designed agents play an important role in helping LLaMA-3-8B outperform the most recent competitive commercial LLMs (i.e., o1, o3-mini, and GPT-4.5) with only 8 billion parameters.
Data Availability
The data used in our work is available at https://github.com/AI-in-Health/RiskAgent
Code Availability
The code that supports the findings of this study is available at https://github.com/AI-in-Health/RiskAgent.

### v2

@@ 4.1 Evaluation on Risk Prediction
We build a benchmark MedRisk consisting of diverse risk problems for evaluating LLMs’ performance for risk prediction. We first collect all available evidence-based tools, i.e., clinical calculators, from clinical-standard source MDCalc (Elovic & Pourmand, 2019). We then review and exclude tools that are not directly predicting disease risks (e.g., those predicting Pregnancy Due Dates and BMI), retaining 387 tools that predict 154 diseases across different scenarios. We use the APIs provided by MDCalc to generate over 15,000 questions with free-text reference answers by requesting GPT-4o (OpenAI, 2023) in Azure to randomly select appropriate tool parameters, 20 quantitative and 20 qualitative questions for each tool. Next, we manually identify and exclude data samples that are unrelated to risk prediction, obtaining 12,352 risk questions across 154 diseases, 86 symptoms, 50 specialties, and 24 organ systems. For robust and efficient evaluation, we convert these questions and free-text references to multiple-choice format by combining the correct answer with three wrong choices. We finally divide the benchmark into MedRisk-Qualitative and -Quantitative, with 6,176 data samples each, according to their question and answer types. For example, the question ‘What is the estimated risk of postoperative pulmonary complications?’ (Answer: High risk) belongs to Qualitative, while the question ‘What is the Risk Score for Asthma Exacerbation (RSE) for this patient?’ (Answer: 30-37%) belongs to Quantitative. Figure 1(b) shows the statistics of the top-30 diseases, top-30 symptoms, top-15 specialties, and top-15 organ systems in the benchmark.
Table 2: Performance of tool learning on the external MEDCALC-BENCH (Khandekar et al., 2025) benchmark, which is categorized into Equation-based (Lab, Physical, Date, Dosage) and Rule-based (Risk, Severity, Diagnosis) tools.
@@
Public LLMs PMC-LLaMA-13B (Wu et al., 2024) 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Meditron-70B (Chen et al., 2023) 0.00 0.00 3.33 0.00 0.00 0.00 3.33 0.38
Mistral-7B (Jiang et al., 2023) 10.09 14.58 1.67 0.00 9.58 7.50 25.00 10.79
LLaMA-3-8B (AI@Meta, 2024) 16.51 25.00 1.67 7.50 11.25 13.75 26.67 16.43
Mixtral-8x7B (Jiang et al., 2024) 22.63 40.00 6.67 17.50 11.25 21.25 15.00 22.35
@@
Commercial LLMs GPT-3.5 (OpenAI, 2022) 20.49 45.00 11.67 17.50 13.33 10.00 31.67 23.69
GPT-4 (OpenAI, 2023) 26.30 71.25 48.33 40.00 27.50 15.00 28.33 37.92
GPT-4o (Hurst et al., 2024) 40.98 82.50 36.67 47.50 32.50 22.50 23.33 46.13
@@
Table 3: Performance of medical reasoning and question answering on three widely used benchmarks MedQA, MedMCQA, and MMLU.
Methods MedQA MedMCQA MMLU
Clinical Knowledge College Medicine Medical Genetics Professional Medicine
MedGemma (Sellergren et al., 2025) 83.6 65.3 80.0 76.9 83.0 68.4
GPT-4o (Hurst et al., 2024) 86.0 77.0 88.3 85.5 96.0 93.8
Claude-3.7-Sonnet (Anthropic, 2024a) 78.0 72.1 85.7 78.0 94.0 89.7
Claude-4-Sonnet (Anthropic, 2024c) 78.6 74.5 86.8 78.6 95.0 91.5
Claude-4-Opus (Anthropic, 2024b) 79.3 77.8 91.7 78.0 96.0 91.9
RiskAgent-GPT-4o (Ours) 87.8 80.8 89.4 86.6 96.0 95.2
Table 4: Ablation study of our RiskAgent, which includes four main components: Tool Library, Decider, Executor, and Reviewer. We take LLaMA-3-8B as the basic LLM and use accuracy to report our results.
@@
Results
In Table 1, we report different methods’ performance on the most common cases across disease, symptom, specialty, and organ system and their overall performance. RiskAgent not only outperforms different types of LLM in all cases, respectively, but also achieves the highest overall accuracy of 78.34% and 76.33% in MedRisk-Qualitative and MedRisk-Quantitative, respectively, doubling the 38.39% accuracy of GPT-4o. It indicates the robustness of our method across different diseases, symptoms, specialties, organ system, and question types, providing a more comprehensive risk prediction solution than previous methods. In detail, compared with large commercial LLMs (e.g., GPT-4o, GPT-o1, and GPT-o3-mini) and open-source LLMs (e.g., Meditron-70B and LLaMA-3-70B), our RiskAgent outperforms them by >>20.0% accuracy, with much (at least 10 times) fewer parameters. This is desirable in resource-limited medical settings. Our method also achieves the best results among AI agents, with accuracy surpassing previous methods by up to 35.83% overall accuracy.
To clearly understand the effectiveness of our approach, we visualize the results on MedRisk-Quantitative in Figure 1(c), which shows that our RiskAgent outperforms previous methods in all cases. These encouraging results demonstrate that our RiskAgent achieves more precise and robust decision support than existing advanced LLMs for complex medical tasks. Importantly, our method uses evidence-based tools for prediction and traces the information source (i.e., evidence) behind our system’s decisions, making the results more transparent and reliable than existing approaches.
4.2 Evaluation on Tool Learning
We further evaluate our method’s tool learning ability on an external benchmark MEDCALC-BENCH (Khandekar et al., 2025), which requires the models to use medical calculations for decision support. Note that our RiskAgent is designed specifically for risk prediction; thus, this external benchmark evaluates our methods’ tool learning ability to new tasks and domains. For evaluation, we follow previous work (Khandekar et al., 2025) to employ zero-shot chain-of-thought (CoT) prompting to report the results. We adopt different backbones, i.e., LLaMA-3-8B (AI@Meta, 2024) and GPT-4o (Hurst et al., 2024), to implement two variants of our method: RiskAgent-8B and RiskAgent-GPT-4o.
Table 2 clearly show that: (i) The RiskAgent-8B model, with only 8B parameters, surpasses all public LLMs, including the larger LLMs Meditron-70B (Chen et al., 2023) and Mixtral-8x7B (Jiang et al., 2024) in major cases. (ii) The RiskAgent-GPT-4o model significantly outperforms all baseline models with an overall accuracy of 67.71%, representing a substantial improvement over the standard GPT-4 and GPT-4o. Most notably, our model exhibits exceptional performance in rule-based diagnosis tasks, achieving 61.67% accuracy in the Diagnosis category - more than double the performance of all baseline methods. Meanwhile, the performance improvements are particularly striking in calculation-heavy categories, with RiskAgent-GPT-4o achieving 97.08% on Physical assessments and 95.00% on Dosage calculations - areas where precise quantitative reasoning is critical for patient safety. It can also be verified by our RiskAgent-8B model, which, outperforms all public and commercial LLMs in dosage calculations.
4.3 Evaluation on Question Answering
We further adopt three widely used benchmarks, MedQA (Jin et al., 2021), MedMCQA (Pal et al., 2022), and MMLU (Hendrycks et al., 2020), to evaluate the performance of our method in general medical reasoning and question answering. As shown in Table 3, RiskAgent achieves the best results on most datasets. This demonstrates that although our method is designed for risk prediction, it exhibits strong generalization capabilities across diverse diagnostic tasks in previously unseen clinical scenarios and maintains robust medical reasoning performance on general questions.
Overall, these encouraging results further underscore the effectiveness and robustness of our approach and its potential to provide reliable decision support across various medical contexts by synergizing language models with validated tools for evidence-based medical recommendations.
5 Analysis
5.1 Ablation Study
@@
Our method introduces four main components, i.e., the Tool Library in the environment, the Decider, the Executor, and the Reviewer, to achieve superior performance. (i) As shown in Table 4, all components contribute to improved performance, proving the effectiveness of each component. (ii) By comparing Basic LLM and Setting (a), which introduces the tool library to enable the basic LLM to use the evidence-based tool for risk prediction, we observe that Setting (a) significantly boosts the performance of the basic LLM by more than 14% accuracy, surpassing larger models like LLaMA-3-70B and GPT-4o. This proves the motivation and effectiveness of leveraging evidence-based tools to achieve improved performance. (iii) Considering that collaborating with tools involves both tool selection and parameter parsing for tool execution, we further report their accuracy. Comparing Setting (a) and Setting (b), we observe that introducing LLM agents can significantly boost performance—improving tool selection accuracy by 9.66%, parameter parsing accuracy by 30.20%, and overall task performance by 18.99%. This proves the effectiveness of our method in separating the collaboration process into tool selection and parameter parsing through the designed Decider and Executor. By distributing responsibilities, our approach reduces the model’s learning burden, enabling accurate and efficient synergy with medical tools. (iv) Finally, we notice that the Reviewer further improves overall performance, showing the effectiveness of reflecting on the analysis process and thereby correcting potential errors.
Figure 3: We evaluate the performance of models on five rare diseases (left) and six types of cancer (right). SM: Systemic Mastocytosis; PM: Primary Myelofibrosis; CML: Chronic Myelogenous Leukemia; IPF: Idiopathic Pulmonary Fibrosis.
Figure 4: The generalization ability of our method: We report the overall accuracy of the basic LLMs (blue bars) and the basic LLMs enhanced using our method (red bars). We evaluate a. different variants of the GPT-4o-series LLMs developed at different times and b. LLaMA-3-series LLMs with varying numbers of model parameters. The polyline and the right y-axis show the improvements in different variants. We can see that the more advanced (a.) and the larger (b.) the basic LLM, the greater the improvements.
5.2 Effect of Rare Diseases and Cancer
We further conduct experiments on rare diseases and cancer to evaluate our method’s robustness in solving diverse complex diseases. For rare (but important) diseases, we follow two clinical-standard rare disease databases, i.e., European Orphanet (https://www.orpha.net/) and the US National Organization for Rare Disorders (NORD) (https://rarediseases.org/), to choose five rare diseases for evaluation. The upper part of Figure 3 shows that RiskAgent achieves the best results. Due to the scarcity of rare disease data, previous LLMs struggle to learn sufficient knowledge during their pretraining, making it difficult for them to provide accurate analysis and predictions for rare diseases. Fortunately, our method collaborates with specialized tools for these rare diseases to leverage their evidence-based knowledge and advantages, achieving significantly improved performance. For cancer, as shown in the bottom part of Figure 3, we can see that our method consistently achieves competitive results with different cancers. It shows the potential of our solution, with only 8B parameters, to provide more accurate and robust risk prediction for complex diseases compared to state-of-the-art LLMs.
5.3 Generalization Analysis
We provide a solution that synergizes LLMs with validated tools to produce accurate, reliable medical results. Thus, our method is agnostic to the underlying LLM, making it adaptable to various LLMs and helping them produce accurate and evidence-based predictions. To this end, we further explore our solution’s generalizability and applicability as LLMs continue to evolve rapidly over time and as the number of parameters increases. We select the GPT-4o-series developed at different times (gpt-4o-2024-05-13, gpt-4o-2024-08-06, and gpt-4o-2024-11-20) and LLaMA-series (LLaMA-3-1B, LLaMA-3-3B, and LLaMA-3-8B), which share the architecture but differ in model size.
[…]
6 Conclusion
In this paper, we present RiskAgent to achieve faithful and evidence-based medical recommendations. By grounding recommendations in valiated tools, RiskAgent offers a resource-efficient, transparent, and trustworthy solution for complex clinical decision-making. Extensive experiments demonstrate that RiskAgent, using only 8 billion parameters, achieves competitive performance with state-of-the-art models in risk prediction, tool learning, and medical reasoning.
The encouraging results highlight the strong potential of our approach across diverse medical domains. For example, instead of extensively fine-tuning LLMs for different tasks, as in common practice, our method collaborates with and leverages validated tools, not only producing trustworthy results but also reducing resource costs, thereby making LLMs more accessible for resource-limited medical applications.
Impact Statement
This work advances clinical AI by providing potentially expert-level risk prediction for resource-limited institutions, as our 8B model runs effectively on accessible hardware. By leveraging validated medical tools, RiskAgent mitigates the hallucination issues common in general LLMs, aligning AI outputs with evidence-based medicine. While this approach enhances transparency, real-world deployment must strictly adhere to privacy regulations regarding patient data. Furthermore, RiskAgent is designed as a decision support tool; clinician oversight remains essential to ensure patient safety and validate tool applicability.
Our study was conducted on retrospective, de-identified data, in which all Protected Health Information (PHI)—such as patient name, sex, gender, and date of birth—was officially de-identified in compliance with the Health Insurance Portability and Accountability Act (HIPAA) standards.